on theses I would like your on aproximation. Thank you so much

1. The squareroot of this problem(2)(7) and
if you would this one square root of (2)(6) this is how they are written in a foil type question

Answer by RAY100(1637)

sqrt2*6=sqrt12=sqrt4*sqrt3=2sqrt3,,,,,(3.464 with calc)
sqrt(2)*(7)=sqrt(14),,,,,,3.742,,,,no factors available
again reducing technique is to multiply given factors, and then find perfect square factors,separate, then take sqrt